• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pkpan

    Log entry time 08:00:27 on June 14, 2009

    Entry number 281261

    keyword=owl shift summary

    On shift: SL: Kai Pan TO: Haijiang Lu 3rd: Haijiang Lu

    Miham presented

    1:00 Beam is back. We follow the "major down time "beam recovery procedure from the "step 5" directly, that is target moved to the Beo , 5uA and without Harp Scan.

    1:05 start #3570/ #22425 (check out run)

    1:10 stop #3570/#22425 because the beam is not in the right position.

    1:10 We ask MCC to tune the beam postion. We start #3571/#22426 (check out run), after the beam moving to the right position.

    1:30 raster size and BPM OK, we stop #3571/#22426, and target moved to He3.

    1:30 NMR 63.16%.

    1:40 VNC trips. Problem seems be fixed.

    1:45 start #3572/#22427 (check out) and find that BPM A(X) is not in the right position. Ask MCC to tune the beam.

    2:15 new Happex Run 31528(stop) and 31529

    2:20 #3573/#22428 (check out)

    2:25 #3574, 3575/#22429, #22430 (check out) with the raster size 10*8 MCC, we found that it hit the corner.

    2:50 #3576, #22431 (check out)

    3:00 #3577, #22432 (check out)

    3:15 a new compton run #17465

    3:15 #3578, #22433 (check out)

    3:30 #3579, #22434 (production run, raster size 5*4 MCC and 4*3.5(mm), 14uA)

    4:02 #3580, #22435 (production run)

    4:25 wire chambers trips again. Disable the alarmed frame,and enable them. But it trips a second time right after . I disabled all of the channels again.

    4:35 beam trips frequently during this time, and MCC called us, saying that there will be a 15 minute beam shut down.

    4:50 beam back

    4:50 start #3581/ #22436 (production run). It seems that wire chamber works now.

    5:21 NMR 63.1%

    5:25 start a new Compton 17466, former one 17465 ending failed

    5:25 start #3582/ #22437 (production run)

    5:52 start #3583/ #22438 (production run)

    6:15 target moved to ref cell (He3 in ref cell now)

    6:22 #3584/ #22439 (calibration run)

    6:48 #3585/ #22440 (calibration run) (Empty ref cell)

    7:15 pump H2 into ref cell

    7:15 #3586/ #22441 (calibration run) (H2 ref cell)

    7:30 a new Compton Run 17467

    7:30 a new Happex Run 31530

    7:30 pump N2 into ref cell

    7:48 #3587/ #22442 (calibration run) (N2 ref cell)

    For the details of calibration run, please refer to Miham's log.